# This Candidate is [MASK]

### v1

@@ B. Correlations
Figure 3 shows how the distribution of prompt-based sentiment scores varies by different covariates. [^28] See appendix table A.1 for detailed descriptives for both sentiment scores and letter length. The job market is known to be quite dynamic. However, there is no evidence of a trend over time in average letter sentiment or its dispersion. There does seem to be an upwards trend in letter length, however.
I find a relatively small statistically significant difference in average sentiment scores ($\sim 3$ pp.) in favor of women. [^29] In contrast, the average length of letters written for male candidates is a bit higher ($\sim 20$ words); but the difference is not statistically significant.
A somewhat surprising finding involves differences across fields of research. Specifically, the letters written for candidates who work in theoretical fields (micro, game theory, mathematical economics) describe them in significantly less favorable terms than candidates in other fields. These kind of differences can have multiple explanations. Of course, one distinct possibility is that the theory applicants were on average a bit less competitive than applicants from other fields. Alternatively, it could be the case that for some reason referees are tougher on theorists. Finally, there could be something in my measure of sentiment that introduces a downward bias in theorists’ letters. [^30] For example, the fact that these letters tend to contain more mathematical expressions might make it harder for the LLM to process them. On the other hand, the letters for econometricians —which do not lack technicalities— score much better..
I do not find statistically significant differences across PhD program rank or region. Finally, candidates with North American or European citizenship tend to be scored a bit higher (anova p-value=0.056).
@@ 3.3 Testing Prompt-based Learning
Table 4: A Test of Prompt-based Learning — Predicting Candidate Sex
Predictive Performance
Precision Recall F1-score Support
Female 0.95 0.98 0.97 479
Male 0.99 0.98 0.99 1419
Accuracy 0.98 1898
Macro Avg. 0.97 0.98 0.98 1898
Weighted Avg. 0.98 0.98 0.98 1898
While it is possible to generate labels to evaluate a sentiment-scoring model, doing so is costly. Moreover, because of confidentiality and other considerations, it is particularly hard to label letters of recommendation.
Nevertheless, it is still possible to test the prompt-based procedure by predicting readily available features of the text. Here I provide an example consisting in predicting the sex of the job market candidate described in the letter.
@@
Table 4 shows the predictive performance for this, admittedly simple, task. Basically, the model gets it right in all but a couple of cases.
@@ 4 Predicting Job Market Outcomes
There is no unambiguously appropriate measure of what constitutes “success” in the academic job market. Job market outcomes are heterogenous on multiple fronts. Consider figure 4, depicting different aspects of candidates’ placement by September after the job market season. Outcomes vary by the sector of employment, the type of position, and whether the job tasks are mainly research, teaching, or non-academic. In addition, employers differ on how they rank according to their research output, with higher ranked employers being markedly more desirable to the average junior academic.
@@
Notes: Policy institutions include central banks, development agencies and think tanks. “Tenure Track” ={Assistant Prof., Lecturer, Senior Lecturer}.
I construct two outcome variables that are meant to encompass, as best as possible, the most desirable positions. I refer to the first outcome variable as “permanent research positions”. They include tenure-track research positions in academia and permanent research positions in policy institutions. The second outcome is the subset of permanent research positions in top-200 ranked universities and top-50 policy institutions. I refer to the latter as “elite positions”.
Figure 5: Dependent Variables
Notes: “Elite” positions are in the top-200 universities and top-50 policy institutions.
4.1 Linear Regression
@@
In order to study the relation between a candidate’s reference letters and their job market outcome, I first look at the regression:
@@
$y_{it}=$ $\>\alpha+\beta_{1}\,\mathrm{avg\_length}_{it}+\beta_{2}\,\mathrm{avg\_sentiment}_{it}+\boldsymbol{X}^{\prime}_{it}\boldsymbol{\gamma}+\lambda_{t}+\epsilon_{it}$
where $y_{it}$ is the job market outcome of candidate $i$, in job market year $t$. The main predictors in these models are the average letter length and average letter polarity. In all regressions, I introduce controls for the candidates’ characteristics (gender, nationality, publications, long PhD), their PhD program’s characteristics (rank, region) and the job market year.
@@
Reference letters are only one input in the recruitment process. Job market outcomes also depend on the candidate’s ability favorably to present themselves in interviews, seminar presentations, and one-on-one meetings with faculty and administrators at potential employers, among other variables. Because these other determinants remain unobserved, the regression estimates must be interpreted with care. Assuming strong candidates get good letters and also tend to perform well in interviews and seminars, one reasonable starting point is to expect regression coefficients to over-estimate the true causal effect that the letters might have on a candidate’s prospects.
Table 5: Dictionary Methods — Predicting Job Market Outcomes
(1) (2) (3) (4)
HGI LM Vader Senticnet
A. Permanent Research Position
Average Letter Length 0.1658∗∗∗ 0.1600∗∗∗ 0.1653∗∗∗ 0.1640∗∗∗
(0.0531) (0.0531) (0.0532) (0.0544)
Average Polarity -0.0251 -0.1110 0.1141 -0.1853
(0.1298) (0.0855) (0.2298) (0.3064)
B. Elite Position
Average Letter Length 0.2780∗∗∗ 0.2771∗∗∗ 0.2766∗∗∗ 0.2807∗∗∗
(0.0513) (0.0515) (0.0514) (0.0521)
Average Letter Polarity 0.0720 -0.0145 0.2628 0.2349
(0.1222) (0.0788) (0.2121) (0.2385)
Candidate’s characteristics Yes
PhD characteristics Yes
Job Market Year Yes
Notes: $N=603$. Robust standard errors in parenthesis. p$<$0.1; ∗∗p$<$0.05; ∗∗∗p$<$0.01. Candidate’s characteristics: sex, nationality, publications, PhD lasting 7 years or longer. PhD program characteristics: rank category, region.
Table 5 presents a summary of my findings when using dictionary methods to measure letter sentiment. Across the board, I find that the lexical method produces sentiment scores that are not predictive of job market outcomes. Notice how, in contrast, average letter length has a strong positive association with the probability of successful placement in the job market. Letters that are on average about 1,000 words longer are associated with a 16–17 pp. higher probability of attaining a permanent research position, and about 28 pp. higher probability of employment in a high-ranked institution.
Table 6: Prompt-based Sentiment — Predicting Job Market Outcomes
Permanent Research Position Elite Position
Average Letter Length 0.1693∗∗∗ 0.2814∗∗∗
(000s words) (0.0524) (0.0509)
Average Prompt-based Polarity 0.5238∗∗∗ 0.5395∗∗∗
(0.1553) (0.1440)
Candidate’s characteristics Yes
PhD characteristics Yes
Job Market Year Yes
Notes: $N=603$. Robust standard errors in parenthesis. p$<$0.1; ∗∗p$<$0.05; ∗∗∗p$<$0.01. Candidate’s characteristics: sex, nationality, publications, PhD lasting 7 years or longer. PhD program characteristics: rank category, region.
In table 6, I present similar regression estimates when using my prompt-based strategy. The results for average letter length are almost identical, which is not too surprising given that letter length and letter polarity are not strongly correlated. On the other hand, in contrast to dictionary methods, the prompt-based sentiment is strongly predictive of good job market outcomes.
@@ B. Elite Position
One point worth emphasizing is how prompt-based learning enables a straightforward interpretation of the regression results, even if they cannot be taken to measure a causal effect. For example, an increase in average polarity of 0.1 means that the LLM assigns a 10 pp. higher probability to filling the prompt with one of the positive sentiment tokens relative to the negative tokens. According to the findings in the table, such as increase is predictive of a 5.2 pp. increase in the expected probability of obtaining a permanent research position, and a 5.4 pp. higher probability of obtaining an “elite” position.
4.2 Nonlinearities
@@
In order to explore nonlinear relationships and interactions, I estimate random forest classifiers with the same outcome and predictor variables. [^31] See Athey and Imbens (2019) for details on random forests. The main hyper-parameters are set based on 5-fold cross-validation. [^32] Table A.2 in the appendix presents the hyper-parameter values. As usual, the estimation of each decision tree is based on a different sample bootstrap and a random subset of the predictors.
@@
I use the fitted random forests to obtain partial dependence curves, which simulate the changes in the predicted probability for each of the outcomes. Formally, the partial dependence curve for the predictors $X_{S}$ is obtained as follows:
@@
where $X_{C}$ are all other predictors and $\mathcal{X}_{S}$ is the support set of $X_{S}$.
In figure 6, I present partial dependence curves for the average letter prompt-based sentiment score. While the random forest confirms the positive association between letter scores and employment prospects, it suggests that the relationship might not be linear.
Figure 6: Partial Dependence Curves — Average Prompt-based Letter Polarity
(a) Permanent Research Position
(b) Elite Position
@@
Note: Partial dependence curves simulate average job market outcomes when the whole sample is assigned (counterfactual) levels of average letter polarity.
The same method can be applied to study interactions between letter sentiment and other predictors. For example, to study heterogeneity across gender lines, I let $X_{S}=\{\mathrm{avg\_sentiment},\mathrm{gender}\}$. The results of these simulations are presented in figure 7.
Figure 7: Interactions — Letter Polarity and Gender
(a) Permanent Research Position
(b) Elite Position
Note: Partial dependence curves simulate average job market outcomes when the whole sample is assigned (counterfactual) values for gender and average letter polarity.
The average predicted probability of obtaining a permanent research position is virtually identical when the sample is assigned either sex, regardless of the letter sentiment. In contrast, women are predicted to have a small advantage in getting hired at top institutions, with the sex gap somewhat narrowing as letters get stronger.
Figure 8: Interactions — Letter Polarity and Field
(a) Permanent Research Position
(b) Elite Position
Note: Partial dependence curves simulate average job market outcomes when the whole sample is assigned (counterfactual) values for field of research and average letter polarity.
Figures 8 and 9 present similar simulations, this time by field of research and PhD rank group respectively.
Figure 9: Interactions — Letter Polarity and PhD Rank
(a) Permanent Research Position
(b) Elite Position
Note: Partial dependence curves simulate average job market outcomes when the whole sample is assigned (counterfactual) values for PhD rank group and average letter polarity.
Clearly, applied fields have a tougher time in the market. Even candidates with very strong letters have distinctly lower predicted probability of attaining a permanent research position or an elite position. In fact, the gap with other fields seems to widen as letters get stronger.
Interestingly, the random forest predictions suggest that the ranking of the PhD program does not make a very sizeable difference in terms of obtaining a research position. On the other hand, elite positions are easier to access for PhD candidates from top ranked universities. The gradient with respect to average letter polarity is similar for all candidates regardless of their PhD rank.
Figure 10: Interactions — Letter Polarity and Letter Length
(a) Permanent Research Position
(b) Elite Position

### v2

@@ 3.2 Reference Letter Sentiment
Figure 2 shows how the distribution of prompt-based sentiment polarity varies by different covariates. I focus on categories where differences are strongly statistically significant. I find a small difference in average sentiment scores in favor of women and against candidates from Asia and from Developing countries.
Letter sentiment tends to reinforce other known signals of higher ability. Candidates from PhD programs in America and from programs in the top 25 have substantially better RLs. The same is true about candidates with top publications, as well as candidates who managed to procure a letter from a top-5 percentile research economist. These statistics are in line with what a general sentiment measure should look like under the general assumption that higher ability candidates get more enthusiastic references. In section 5 below, I further analyze the determinants of sentiment polarity as well as other RL features extracted via prompts.
@@ 3.3 Testing Prompt-based Learning
While it is possible to generate labels to evaluate a sentiment-scoring model, doing so is costly and not necessarily convincing since it would be hard to hire reviewers with subjective opinions similar to those of hiring committees at research economics departments. Moreover, because of confidentiality and other considerations, it is particularly hard to label RLs.
Nevertheless, it is still possible to test the prompt-based procedure by predicting readily available meta-data that one can safely assume is contained in most RLs. Here I provide an example consisting in predicting the sex and research field of the job market candidate described in the letter. For this purpose, I addressed the model with the prompts and verbalizers detailed in table 4.
Table 4: Tests of Prompt-based Learning
Test Prompt ($\mathcal{T}$) Verbalizer ($\mathcal{V}$)
@@
[…]
Table 5: Prompt-based Learning Performance
A. Candidate Sex Precision Recall F1-score Support
Female 1.00 0.99 1.00 507
Male 1.00 1.00 1.00 1461
Accuracy 1.00 1968
Macro Avg. 1.00 1.00 1.00 1968
Weighted Avg. 1.00 1.00 1.00 1968
B. Field of Research Precision Recall F1-score Support
Applied 0.83 0.82 0.82 673
Econometrics 0.73 0.85 0.79 108
Finance 0.94 0.87 0.90 532
Macroeconomics 0.79 0.85 0.82 466
Theory 0.75 0.74 0.74 189
Accuracy 0.83 1968
Macro Avg. 0.81 0.83 0.82 1968
Weighted Avg. 0.84 0.83 0.83 1968
Table 5 shows the model’s predictive performance for these tasks. The sex test is admittedly simple and could probably be solved by a dictionary counting pronouns as well. Basically, prompt-based learning gets it right in all but four cases.
Predicting field of research is not an easy task and it would probably have high error rates even if a human read and labeled each RL. First, not every RL really mentions the field of research of the candidate. There are “teaching” references and other short letters that do not really delve into the candidate’s research. Second, lots of candidates do work on multiple fields, or fields that are not pure cases (e.g. ‘applied theory’, ’central banking’, ‘macro labor’, ‘structural econometrics’, etc.). In sum, classifying candidates into fields based on a RL is far from trivial. Given these considerations, prompt-based learning performs remarkably well, with accuracy and average F-1 scores well above 80%.
@@ 4 Predicting Job Market Outcomes
There are multiple potentially appropriate measures of what constitutes “success” in the academic job market. Job market outcomes are heterogenous on multiple fronts. Consider figure 3, depicting different aspects of candidates’ placement by September after the job market season. Outcomes vary by the sector of employment and the type of position. In addition, employers differ on how they rank according to their research output, with higher ranked employers being markedly more desirable to the average junior academic. [^31] In order to determine the employer’s ranking, I once again use Tilburg’s ranking for academic institutions. For policy institutions, I use the econphd.net ranking (https://econphd.econwiki.com/rank/rallec.htm), which focuses on publications and has a transparent methodology.
@@
Notes: NN=645. Policy institutions include central banks, development agencies and think tanks.
In what follows, I define my baseline “success” outcome as attaining a top-100 academic or top-50 policy job. For some clarity on these cutoffs, consider that economics departments ranked 98–102 in the Tilburg ranking are NHH (Norwegian School of Economics), INSEAD, the California Institute of Technology and the University of Gothenburg. Institutions ranked in the top-50 of the policy ranking include the Federal Reserve Banks and the Board of Governors, as well as the ECB and a few other major European central banks, plus the World Bank, the IMF and a couple other major development institutes. In figure 3, the baseline cutoffs are depicted by vertical red lines. Clearly, only a minority of candidates in my sample attained success based on this measure.
As a robustness check, I also define four alternative outcome variables. The main modification in alternative 1 is to exclude post-docs from the definition of success, which arguably are less desirable than tenure-track positions even if they are sponsored by top universities. Alternative 2 considers the possibility that some positions at tech companies are as desirable as top academic placements. Finally, alternatives 3 and 4 simple change the cutoffs for success. The full definitions of these variables, as well as basic descriptive statistics, can be found in table 6.
Table 6: Job Market Success: baseline and alternative definitions
Outcome Definition $\mathbf{N}_{success}$ $\mathbf{\%}_{success}$
Baseline Top-100 Academic Employers & Top-50 Policy Institutions 175 27.13
Alt. 1 Exclude teaching positions (liberal arts colleges) and post-docs from the success group 115 17.83
Alt. 2 Add some highly paid tech jobs (Amazon, Meta, etc.) to the success group 190 29.46
Alt. 3 Use somewhat more forgiving cutoffs: top-150 academic and top-75 policy 243 37.67
Alt. 4 Use somewhat more restrictive cutoffs: top-75 academic and top-35 policy 150 23.26
Notes: NN=645.
Figure 4: Average RL Sentiment and Success in the Academic Job Market
Notes: NN=645. $N_{success}$ shows by how many alternative definitions the candidate had success in the job market.
Figure 4 depicts the relationship between average RL quality and success in the academic job market. It is not hard to see that candidates who succeeded according to most definitions are clustered above and to the right relative to those who did not succeed.
4.1 Linear Regressions
@@
In order to quantify the relationship between the sentiment contents of a candidate’s RLs and their job market outcome, I first look at the regression model:
@@
$y_{it}=$ $\>\alpha+\beta_{1}\,\mathrm{avg\_RL\_length}_{it}+\beta_{2}\,\mathrm{avg\_RL\_sentiment}_{it}+\boldsymbol{X}^{\prime}_{it}\boldsymbol{\gamma}+\lambda_{t}+\epsilon_{it}$ (3)
where $y_{it}$ is the job market outcome of candidate $i$, in job market year $t$. The main predictors in these models are the average RL length (in thousands of words) and average RL sentiment polarity (in p.p.). [^32] In the interest of full clarity, the regressors are the average RL word count divided by 1,000 and the average sentiment polarity score multiplied by 100. These adjustments ease the interpretation of the coefficients by eliminating unnecessary decimal points. All regressions include controls for the number of RLs submitted on behalf of the candidate. Depending on the exact specification, the vector of control variables may also include: the candidate’s individual characteristics (sex, region of origin, field of research, publications, long PhD), the PhD program’s characteristics (rank, region), the job market year, and whether the letter writers are senior (specifically, the number of top-5 writers).
@@
RLs are only one input in the recruitment process. Job market outcomes also depend on how the candidates present themselves in interviews, seminar presentations, and one-on-one meetings with faculty and administrators at potential employers, among other variables. Because these other determinants remain unobserved, the regression estimates cannot be interpreted causally. [^33] Assuming strong candidates get good letters and also tend to perform well in interviews and seminars, one reasonable expectation would be for regression coefficients to over-estimate the true causal effect that RL sentiment might have on a candidate’s probability of success.
However, in light of the theoretical discussion regarding whether referees can be trusted to offer honest assessments and whether hiring committees are properly incentivized to select the best candidates, it is still interesting to determine whether a statistically significant association exists between good letters and good job market outcomes. In this section and the next, I show that not only do higher average sentiment RLs are predictive of better outcomes, but also that the degree of disagreement among the letter writers—sentiment dispersion—is negatively associated with success in the market.
Table 7: Predicting Success in the Academic Job Market
Baseline Outcome: top-100 academic and top-50 policy jobs
(1) (2) (3) (4) (5)
Average Letter Length 0.2036∗∗∗ 0.2510∗∗∗ 0.2371∗∗∗ 0.2416∗∗∗ 0.2447∗∗∗
(000s words) (0.0473) (0.0469) (0.0470) (0.0503) (0.0506)
[0.0495] [0.0496] [0.0501] [0.0548] [0.0546]
||0.0470|| ||0.0450|| ||0.0446|| ||0.0501|| ||0.0515||
Average Letter Sentiment 0.0304∗∗∗ 0.0276∗∗∗ 0.0230∗∗∗ 0.0233∗∗∗ 0.0227∗∗∗
($p.$) (0.0071) (0.0067) (0.0069) (0.0070) (0.0070)
[0.0067] [0.0062] [0.0063] [0.0064] [0.0064]
[…]
Number of Letters Yes Yes Yes Yes Yes
Candidate’s characteristics No Yes Yes Yes Yes
PhD characteristics No No Yes Yes Yes
Job Market Year No No No Yes Yes
Number of top-5% Writers No No No No Yes
Adjusted $R^{2}$ 0.0894 0.1346 0.1408 0.1323 0.1303
[…]
Table 8: Alternative Definitions of Success
[…]
Average Letter Length 0.2447∗∗∗ 0.1950∗∗∗ 0.2180∗∗∗ 0.2578∗∗∗ 0.1887∗∗∗
[…]
[0.0064]∗∗∗ [0.0057]∗∗∗ [0.0068]∗∗∗ [0.0081]∗∗∗ [0.0064]∗∗∗
[…]
Number of Letters Yes Yes Yes Yes Yes
Candidate’s characteristics Yes Yes Yes Yes Yes
PhD characteristics Yes Yes Yes Yes Yes
Job Market Year Yes Yes Yes Yes Yes
Number of top-5% Writers Yes Yes Yes Yes Yes
Adjusted $R^{2}$ 0.1303 0.1308 0.1175 0.1392 0.1125
[…]
Table 7 shows regression estimates of equation (3) for the baseline outcome and with different sets of control variables. [^34] Table A.2 in the appendix presents the coefficient estimates for the control variables. Regardless of the specifics, both RL length and sentiment have a strong and highly statistically significant relationship with the probability of success. [^35] Throughout the paper I present standard errors, and the corresponding levels of significance, for three different assumptions regarding the joint distribution of the error terms in equation (3). HC0 standard errors assume the $\epsilon_{it}$ are independently distributed but heteroskedastic. Clustered standard errors allow for dependence across errors either within PhD programs (due to norms regarding how to write RLs, for example) or within clusters defined by the market year and the PhD rank group (due to competition among candidates for similar jobs). Interestingly, all the evidence suggests that the length of the RLs acts as an implicit positive signal in and of itself. A one-thousand word delta in average RL length is predictive of a very sizeable increase in the expected probability of success.
@@
Prompt-based sentiment extraction enables a straightforward interpretation of the regression results, even if they cannot be taken to measure a causal effect. For example, an increase in average polarity of 0.1 means that the LLM assigns a 10 pp. higher probability to filling the prompt with one of the positive sentiment tokens relative to the negative tokens. According to the findings in the table, such as increase is predictive of a much higher ($\sim$23 p.p.) expected probability of obtaining a position in a top research job either in academia or in a policy institution.
In table 8, I present similar estimates for the different definitions of success in the academic job market. The results are qualitatively identical in all cases. Among the different dependent variables, alternative 1 (removing post-docs and teaching jobs from the definition of success) seems to have a somewhat weaker dependence on RL sentiment.
4.2 Sentiment Dispersion
Almost without exception, employers in the job market for academic economists require interested applicants to submit multiple confidential RLs. This convention suggests that employers are more than willing to tradeoff the extra time required to assess the additional information for insurance against the risk of biased signals from poorly incentivized letter writers.
An open question is whether disagreement among the many letter writers is reflected in worse job market outcomes for the candidates. In order to study this issue, I focus on a sub-sample of “complete applications,” i.e. candidates for whom I have at least three RLs including the letter from the main research adviser. [^36] Table A.3 in the appendix replicates the analysis in table 8 for this sub-sample. The results are practically identical. Figure A.1 in the appendix shows that the distributions of sentiment polarity scores for different letter writers are almost identical. [^37] In unreported analysis, I found that the adviser’s RL sentiment and other writer’s RL sentiment have undistinguishable effects on job market outcomes. However, within-candidate variation in RL sentiment is quite high, accounting for over 45% of the total variance of sentiment. [^38] An alternative way to illustrate this point is by looking at correlations across letter writer scores. Excluding the fourth letter, which is available only for a minority of candidates, the correlations range between 0.27 and 0.40. In other words, letter writers tend to be aligned but are far from unanimous.
Table 9: Sentiment Dispersion: summary stats
Mean SD Min Max
Sentiment Average 0.0722 0.0247 0.0013 0.1579
Sentiment Range 0.0467 0.0284 0.0027 0.1505
Sentiment Mean Absolute Deviation 0.0172 0.0104 0.0010 0.0583
Sentiment Standard Deviation 0.0233 0.0141 0.0014 0.0770
Notes: NN=553 (complete applications sub-sample). The range is the difference between the max letter writer score and the min letter writer score for each candidate.
Table 9 shows summary statistics for alternative ways to measure sentiment dispersion among letter writers. The sentiment range—the measure that I will use as a baseline—is defined as the difference in sentiment scores between the writer with the highest score and the writer with the lowest score. For robustness, I will also consider the mean absolute deviation (MAD) and the standard deviation (SD) of scores across letter writers. As is clear from the table, all measures of dispersion are quite high relative to the average. They also vary significantly across candidates.
Figure 5: Sentiment Dispersion and Success in the Academic Job Market
Notes: NN=553 (complete application sub-sample). $N_{success}$ shows by how many alternative definitions the candidate had success in the job market.
Figure 5 depicts the relationship between average sentiment, sentiment dispersion and success in the academic job market. Candidates who succeeded according to most definitions are clustered above and to the left relative to those who did not succeed.
To confirm this observation, I estimate the following regression:
$y_{it}=$ $\ \>\alpha+\beta_{1}\,\mathrm{avg\_RL\_length}_{it}+\beta_{2}\,\mathrm{avg\_RL\_sentiment}_{it}+\beta_{3}\,\mathrm{sentiment\_dispersion}_{it}$ (4)
$\>+\boldsymbol{X}^{\prime}_{it}\boldsymbol{\gamma}+\lambda_{t}+\epsilon_{it}$
where $\mathrm{sentiment\_dispersion}_{it}$ is one of the measures of disagreement across letter writers. Table 10 confirms that, regardless of the outcome definition, sentiment range has a negative relationship to success in the job market. Table A.4 in the appendix shows similar results for alternative measures of sentiment dispersion. In summary, these estimates suggest that RL sentiment has a determinant effect on job market outcomes and that hiring committees fully exploit the information contained in RLs.
Table 10: Sentiment Dispersion and Job Market Success
[…]
Average Letter Sentiment 0.0368∗∗∗ 0.0245∗∗∗ 0.0381∗∗∗ 0.0403∗∗∗ 0.0350∗∗∗
[…]
Four Letters & Avg. Letter Length Yes Yes Yes Yes Yes
Candidate’s characteristics Yes Yes Yes Yes Yes
PhD characteristics Yes Yes Yes Yes Yes
Job Market Year Yes Yes Yes Yes Yes
Number of top-5% Writers Yes Yes Yes Yes Yes
Adjusted $R^{2}$ 0.1324 0.1397 0.1348 0.1389 0.1072
[…]
4.3 Random Forest
@@
In order to explore nonlinear relationships and interactions, I estimate a random forest classifier with the baseline outcome as target and the same predictor variables. [^39] See Athey and Imbens (2019) for details on random forests. The main hyper-parameters are set based on 5-fold cross-validation. [^40] Table A.5 in the appendix presents the hyper-parameter values. As usual, the training of each decision tree is based on a different sample bootstrap and a random subset of the predictors. I apply inverse probability weighting to address the imbalance in the outcome variable. The random forest is trained on the complete applications sub-sample.
One advantage of random forests is that they provide an easy way to assess the predictive performance of the model without the need to set aside a holdout sample. Specifically, it is possible to evaluate each tree in the forest by assessing its predictive performance for the sample elements that were not included in the bootstrap sample used to train the tree. This is referred to as ‘out-of-bag’ (OOB) performance.
Table 11: Random Forest Out-of-Bag Performance
Baseline Outcome Precision Recall F1-score Support
Failure 0.80 0.75 0.77 402
Success 0.42 0.48 0.45 151
Accuracy 0.68 553
Macro Avg. 0.61 0.62 0.61 553
Weighted Avg. 0.69 0.68 0.69 553
Table 11 presents the OOB performance metrics for the prediction of the baseline outcome. Clearly, predicting job market outcomes is much harder than predicting text meta-data. Regardless, the overall accuracy score of 0.68 is not horrible for a random forest trained on 553 examples.
Figure 6: Random Forest Permutation Importances
Notes: The box plots depict the accuracy decrease distributions for 30 independent re-shuffles of each predictive feature.
Figure 6 presents the relative importance of different predictive features. Specifically, I use permutation importances, which measure the average decrease in the OOB accuracy score when the corresponding feature is shuffled. [^41] Specifically, the feature is re-shuffled and the random forest re-trained 30 times. The average RL sentiment score is the most important predictive feature, with a median decrease in accuracy above 6pp. Average RL length and the sentiment range come second and third, respectively.
@@
I use the trained random forest to obtain partial dependence curves, which simulate the changes in the predicted probability of the outcome for counterfactual values of each feature. Formally, the partial dependence curve for the predictor $X_{S}$ is obtained as follows:
@@
where $X_{C}$ are all other predictors and $\mathcal{X}_{S}$ is the support set of $X_{S}$. It is important to emphasize that the $\mathrm{pd}$ curves try to approximate the conditional expectation function without assuming linearity. They do not provide causal estimates, which can only emerge from an experiment.
In figure 7, I present $\mathrm{pd}$ curves for the main predictors: average RL length, average RL sentiment, and sentiment range. For comparison, I also plot the corresponding simulation based on a linear regression. [^42] The regression in the plot is not exactly the same as the regression in the first column of table 10 because I use inverse probability weights in order to match the random forest fit.
While the random forest confirms the positive associations between RL length, RL sentiment scores and employment outcomes, it suggests that the relationship might not be linear. In particular, the curves seem to have steep jumps after crossing certain thresholds. The random forest also point to the relationship between RL sentiment range and job market success being less pronounced than the regression coefficients suggest.
Figure 7: Random Forest Partial Dependence Curves
@@
Notes: Partial dependence curves simulate average job market outcomes when the whole sample is assigned counterfactual levels of the predictor. The vertical marks on the x-axis show deciles of the predictor’s distribution.
The same method can be applied to study interactions between predictors. Figure 8 (left) shows the average probabilities of success when both RL length and sentiment are assigned counterfactual levels. Clearly, these two signals positively interact, with the lowest probability of success in the lower left corner and the highest probability of success in the upper right corner.
The plot on the right of the figure shows simulations for Avg. RL sentiment and RL sentiment range. In this case, the highest probability of success is predicted for candidates with high average RL sentiment and very low sentiment range. Conversely, the lower right corner of the plot has the lowest predicted probability of success.
Figure 8: Partial Dependence Contours